### v2

# Looks Too Good To Be True: An Information-Theoretic Analysis of Hallucinations in Generative Restoration Models

Abstract
The pursuit of high perceptual quality in image restoration has driven the development of revolutionary generative models, capable of producing results often visually indistinguishable from real data. However, as their perceptual quality continues to improve, these models also exhibit a growing tendency to generate hallucinations – realistic-looking details that do not exist in the ground truth images. The presence of hallucinations introduces uncertainty regarding the reliability of the models’ predictions, raising major concerns about their practical application. In this paper, we employ information-theory tools to investigate this phenomenon, revealing a fundamental tradeoff between uncertainty and perception. We rigorously analyze the relationship between these two factors, proving that the global minimal uncertainty in generative models grows in tandem with perception. In particular, we define the inherent uncertainty of the restoration problem and show that attaining perfect perceptual quality entails at least twice this uncertainty. Additionally, we establish a relation between mean squared-error distortion, uncertainty and perception, through which we prove the aforementioned uncertainly-perception tradeoff induces the well-known perception-distortion tradeoff. This work uncovers fundamental limitations of generative models in achieving both high perceptual quality and reliable predictions for image restoration. We demonstrate our theoretical findings through an analysis of single image super-resolution algorithms. Our work aims to raise awareness among practitioners about this inherent tradeoff, empowering them to make informed decisions and potentially prioritize safety over perceptual performance.
1 Introduction
Restoration tasks and inverse problems impact many scientific and engineering disciplines, as well as healthcare, education, communication and art. Generative artificial intelligence has transformed the field of inverse problems due to its unprecedented ability to infer missing information and restore corrupted data. In the realm of image restoration, the quest for high perceptual quality has led to a new generation of generative models  [33, 65, 18, 17, 70], capable of producing outputs of remarkable realism, virtually indistinguishable from true images.
While powerful, growing empirical evidence indicates that generative models are susceptible to hallucinations [30], characterized by the generation of seemingly authentic content that deviates from the original input data, hindering applications where faithfulness is crucial. The root cause of hallucination lies in the ill-posed nature of restoration problems, where multiple possible solutions can explain the observed measurements, leading to uncertainty in the estimation process.
Concerns surrounding hallucinations have prompted the development of uncertainty quantification methods, designed to evaluate the reliability of generated outputs. These approaches offer crucial insights into the model’s confidence in its predictions, empowering users to assess potential deviations from the original data and make informed decisions. Despite these advancements, the intricate relationship between achieving high perceptual quality and the extent of uncertainty remains an understudied area.
This paper establishes the theoretical relationship between uncertainty and perception, demonstrating through rigorous analysis that the global minimal uncertainty in generative models increases with the level of desired perceptual quality (see illustration in Figure 1). Leveraging information theory, we quantify uncertainty using the entropy of the recovery error [19], while we measure perceptual quality via conditional divergence between the distributions of the true and recovered images [51]. Our main contribution are as follows:
  1. 1.
    We introduce a definition for the inherent uncertainty $U_{\text{Inherent}}$ of an inverse problem, and formulate the uncertainty-perception (UP) function, seeking the minimal attainable uncertainty for a given perceptual index. We prove the UP function is globally lower-bounded by $U_{\text{Inherent}}$ (Theorem 1).
  2. 2.
    We prove a fundamental trade-off between uncertainty and perception under any underlying data distribution, restoration problem or model (Theorem 1). Specifically, the entropy power of the recovery error exhibits a lower bound inversely related to the Rényi divergence between the true and recovered image distributions (Theorem 3). This shows that perfect perceptual quality requires at least twice the inherent uncertainty $U_{\text{Inherent}}$.
  3. 3.
    We establish a relationship between uncertainty and mean squared error (MSE) distortion, demonstrating that the uncertainty-perception trade-off induces the well-known distortion-perception trade-off (Theorem 4). We empirically validate all theoretical findings through experiments with image super-resolution algorithms (Section 4.3).
We aim to provide practitioners with a deeper understanding of the tradeoff between uncertainty and perceptual quality, allowing them to strategically navigate this balance and prioritize safety when deploying generative models in real-world, sensitive applications.
Figure 1: Illustration of Theorem 3. In inverse problems, the minimal attainable uncertainty exhibits a lower bound that begins at the inherent uncertainty $U_{\text{Inherent}}$ of the problem (Definition 2) and increases up to twice this value as we approach perfect perceptual quality of the recovery.
2 Related Work
Recent work in image restoration has made significant strides in both perceptual quality assessment and uncertainty quantification, largely independently. Below, we outline the main trends in research on these topics, laying the foundation for our framework.
Perception Quantification   Perceptual quality in restoration tasks encompasses how humans perceive the output, considering visual fidelity, similarity to the original, and absence of artifacts. While traditional metrics like PSNR and SSIM [72] capture basic similarity, they miss finer details and higher-level structures. Learned metrics like LPIPS [75], VGG-loss [64], and DISTS [23] offer improvements but still operate on pixel or patch level, potentially overlooking holistic aspects. Recently, researchers have leveraged image-level embeddings from large vision models like DINO [15] and CLIP [54] to capture high-level similarity. Here, we follow previous works [51, 12, 31] and adopt a mathematical notion of perceptual quality defined as the divergence between probability densities.
Uncertainty Quantification   Uncertainty quantification techniques can be broadly categorized into two main paradigms: Bayesian estimation and frequentist approaches. The Bayesian paradigm defines uncertainty by assuming a distribution over the model parameters and/or activation functions [1]. The most prevalent approach is Bayesian neural networks [47, 68, 35], which are stochastic models trained using Bayesian inference. To improve efficiency, approximation methods have been developed, including Monte Carlo dropout [25, 26], stochastic gradient Markov chain Monte Carlo [59, 16], Laplacian approximations [55] and variational inference [14, 46, 53]. Alternative Bayesian techniques encompass deep Gaussian processes [20], deep ensembles [7, 34], and deep Bayesian active learning [27]. In contrast to Bayesian methods, frequentist approaches operate assume fixed model parameters with no underlying distribution. Examples of such distribution-free techniques are model ensembles [41, 52], bootstrap [36, 2], interval regression [52, 37, 73] and quantile regression [28, 56].
An emerging approach in recent years is conformal prediction [3, 62], which leverages a labeled calibration dataset to convert point estimates into prediction regions. Conformal methods require no retraining, computationally efficient, and provide coverage guarantees in finite samples [44]. These works include conformalized quantile regression [56, 61, 6], conformal risk control [5, 8, 4], and semantic uncertainty intervals for generative adversarial networks [60]. The authors of [40] introduce the notion of conformal prediction masks, interpretable image masks with rigorous statistical guarantees for image restoration, highlighting regions of high uncertainty in the recovered images. Please see [66] for an extensive survey of distribution-free conformal prediction methods. A recent approach [10] introduces a principal uncertainty quantification method for image restoration that considers spatial relationships within the image to derive uncertainty intervals that are guaranteed to include the true unseen image with a user-defined confidence probabilities. While the above studies offer a variety of approaches for quantifying uncertainty, a rigours analysis of the relationship between uncertainty and perception remains underexplored in the context of image restoration.
The Distortion-Perception Tradeoff   The most relevant studies to our research are the work on the distortion-uncertainty tradeoff [12] and its follow-ups [24, 13, 11]. A key finding in [12] establishes a convex tradeoff between perceptual quality and distortion in image restoration, applicable to any distortion measure and distribution. Moreover, perfect perceptual quality comes at the expense of no more than 3dB in PSNR. The work in [24] extends this, providing closed-form expressions for the tradeoff when MSE distortion and Wasserstein-2 distance are considered as distortion and perception measures respectively. In [51], it is shown that the Lipschitz constant of any deterministic estimator grows to infinity as it approaches perfect perception. Our work centers on uncertainty, reflecting the confidence level of a recovery algorithm. Thus, our works complements these studies by introducing uncertainty into the equation and exploring its impact on the perception-distortion relationship.
3 Problem Formulation
We adopt a Bayesian perspective to address inverse problems, wherein we seek to recover a random vector $X\in\mathbb{R}^{d}$ from its observations, represented by another random vector $Y={\mathcal{M}}(X)\in\mathbb{R}^{d^{\prime}}$. Here ${\mathcal{M}}:{\mathbb{R}}^{d}\rightarrow{\mathbb{R}}^{d^{\prime}}$ is a non-invertible degradation function, implying $X$ cannot be perfectly recovered from $Y$. Formally:
Definition 1.
A degradation function ${\mathcal{M}}$ said to be invariable if, the conditional probability $p_{X|Y}(\cdot|y)$ is a Dirac delta function for almost every $y$ in the support of the distribution $p_{Y}$ of $Y$.
The restoration process involves constructing a estimator $\hat{X}\in\mathbb{R}^{d}$ to estimate $X$ from $Y$, inducing conditional probability $p_{\hat{X}|Y}$. The estimation process forms a Markov chain $X\to Y\to\hat{X}$, implying that $X$ and $\hat{X}$ are statistically independent given $Y$.
In this paper, we analyze estimators $\hat{X}$ with respect to two performance criteria: perception and uncertainty. To assess perceptual quality, we follow a theoretical approach, similar to previous works [74, 12], and measure perception using conditional divergence between $X$ and $\hat{X}$ defined as
$D_{v}(X,\hat{X}\big{|}Y)\triangleq\mathbb{E}_{y\sim p_{Y}}\left[D_{v}\big{(}p_{X|Y=y},p_{\hat{X}|Y=y}\big{)}\right],$ (1)
where $D_{v}$ stands for general divergence function. When an estimator attains a low value of the metric above, we say it exhibits high perceptual quality. When it comes to uncertainty, there are diverse practical methods to quantify it [29, 1]. However, for our analysis, we aim to identify a fundamental understanding of uncertainty. Therefore, we adopt the concept of entropy power from information theory, which assesses the statistical spread of a random variable. For the definition of entropy power and other relevant background, we refer the reader to Appendix A. Utilizing entropy power, we formally define the inherent uncertainty intrinsic to the restoration problem as follows
Definition 2.
The inherent uncertainty in estimating $X$ from $Y$ is defined as:
$U_{\text{Inherent}}\triangleq N(X|Y)=\frac{1}{2\pi e}e^{\frac{2}{d}h(X|Y)},$
where $h(X|Y)$ denotes the entropy of $X$ given $Y$.
The inherent uncertainty quantifies the information irrevocably lost during observation, acting as a fundamental limit on the recovery of $X$ from $Y$, regardless of the estimation method. Notably, when the degradation process is invertible, this inherent uncertainty becomes zero $U_{\text{Inherent}}=0$, reflecting the possibility of perfect recovery of $X$ with complete confidence.
We now turn our attention to the main focus of this paper, the uncertainty-perception (UP) function:
$U(P)\triangleq\underset{p_{\hat{X}|Y}}{\min}\;\Big{\{}N(\hat{X}-X|Y)\;:\;D_{v}(X,\hat{X}\big{|}Y)\leq P\Big{\}}.$ (2)
In essence, $U(P)$ represents the minimum uncertainty achievable by an estimator with perception quality of at least $P$, given the side information within the observation $Y$. In contrast to the perception-distortion function [12], the above objective prioritizes the information content of error signals over their mere energy, and its minimization promotes concentrated errors for robust and reliable predictions. The following example offers intuition into the typical behavior of this function.
Example 1.
Consider $Y=X+W$ where $X\sim{\mathcal{N}}(0,1)$ and $W\sim{\mathcal{N}}(0,\sigma^{2})$ are independent. Let the perception measure be the symmetric Kullback–Leibler (KL) divergence $D_{\text{SKL}}$ and assume stochastic estimators of the form $\hat{X}=\mathbb{E}\left[X|Y\right]+Z$ where $Z\sim{\mathcal{N}}(0,\sigma_{z}^{2})$ is independent of $Y$. As derived in Appendix B, the UP function admits a closed form expression in this case, given by
$U(P)=N(X|Y)\Big{[}1+\left(P+1-\sqrt{(P+1)^{2}-1}\right)^{2}\Big{]},\text{ where }N(X|Y)=\sigma^{2}/(1+\sigma^{2}).$
The above result, illustrated in Appendix B, demonstrates the minimal attainable uncertainty increases as the perception quality improves. Moreover, The above example suggests a structure for uncertainty-perception function $U(P)$, which fundamentally relies on the inherent uncertainty $N(X|Y)$. Remarkably, the following section shows that this dependency generalizes beyond the specific example presented here, where its particular form is determined by the underlying distributions, along with the specific perception measure employed.
Remark   One may consider the following alternative formulation
$\tilde{U}(P)\triangleq\underset{p_{\hat{X}|Y}}{\min}\;\Big{\{}N(\hat{X}-X)\;:\;D_{v}(X,\hat{X}\big{|}Y)\leq P\Big{\}}.$ (3)
The alternative objective quantifies uncertainty as the entropy power of the error, independent of the side information $Y$. While potentially insightful, this approach may overestimate uncertainty since $N(\hat{X}-X|Y)\leq N(\hat{X}-X)$ where equality holds if and only if the error ${\mathcal{E}}=\hat{X}-X$ is independent of $Y$. Although further investigation is warranted, we hypothesize that the behavior of function (3) mirrors that of the UP function (2), which we examine in detail in the following section.
4 The Uncertainty-Perception Tradeoff
Thus far, we have formulated the uncertainty-perception function and elucidated its underlying rationale. We now proceed to derive its key properties, including a detailed analysis for the case where Rényi divergence serves as the measure of perceptual quality. Subsequently, we establish a direct link between the UP function and the well-known distortion-perception tradeoff. Finally, we demonstrate our theoretical findings through experiments on image super-resolution.
4.1 The Uncertainty-Perception Plane
The following theorem establishes general properties of the uncertainty-perception function, $U(P)$, irrespective of the specific distributions and divergence measures chosen.
Theorem 1.
The uncertainty-perception function $U(P)$ displays the following properties
  1. 1.
    Quasi-linearity (monotonically non-increasing and continuous):
    $\min\Big{(}U(P_{1}),U(P_{2})\Big{)}\leq U\Big{(}\lambda P_{1}+(1-\lambda)P_{2}\Big{)}\leq\max\Big{(}U(P_{1}),U(P_{2})\Big{)},\;\forall\lambda\in[0,1]$
  2. 2.
    Boundlessness:
    $N(X|Y)\leq U(P)\leq 2N(X_{G}|Y),$
where $X_{G}$ is a zero-mean Gaussian random variable with covariance identical to $X$. The inherent uncertainty is upper bounded by $N(X_{G}|Y)$, which depends on the deviation of $X$ from Gaussianity.
The theorem establishes a fundamental tradeoff between perceptual quality and uncertainty in image restoration, regardless of the specific divergence measure, data distributions, or restoration model employed. This tradeoff is fundamentally linked to the inherent uncertainty $N(X|Y)$ arising from the information loss during the observation process. Notably, the upper bound of $N(X|Y)$ can be expressed as
$N(X_{G}|Y)=N(X|Y)e^{\frac{2}{d}D_{KL}(X,X_{G}|Y)}.$ (4)
This shows that as $X$ approaches Gaussianity, $N(X|Y)$ approaches $N(X_{G}|Y)$. However, concurrently, it implies in general higher values of $N(X|Y)$ due to Lemma 1 of Appendix A. This finding yields a surprising insight: for multivariate Gaussian distributions, perfect perceptual quality comes at the expense of exactly twice the inherent uncertainty of the problem.
The next theorem shows that for a fixed perceptual index $P$, the optimal algorithms lie on the boundary of the constraint set. This facilitates the optimization task, as it restricts the search space to the boundary points.
Theorem 2.
Assume $D_{v}(X,\hat{X}\big{|}Y)$ is convex in its second argument. Then, for any $P\geq 0$, the minimum is attained on the boundary where $D_{v}(X,\hat{X}\big{|}Y)=P$.
While the above theorems describe important characteristics of the uncertainty-perception function, additional assumptions are needed to gain deeper insights. Therefore, we now focus on Rényi divergence as our perception measure. Rényi divergence is a versatile family of divergence functions parameterized by an order $0\leq r$, encompassing the well-known KL divergence as a special case when $r=1$. This divergence plays a critical role in in analyzing Bayesian estimators and numerous information theory calculations [69]. Importantly, it is also closely related to other distance metrics used in probability and statistics, such as the Wasserstein and Hellinger distances. Focusing on the case where $r=1/2$, we arrive at:
$U(P)=\underset{p_{\hat{X}|Y}}{\min}\Big{\{}N(\hat{X}-X|Y)\;:\;D_{1/2}(X,\hat{X}\big{|}Y)\leq P\Big{\}}.$ (5)
While we set $r=1/2$ to facilitate our derivations, it is important to note that all orders $r\in(0,1)$ are equivalent (see Appendix A). Consequently, given this equivalence and the close relationship between Rényi divergence and other metrics, analyzing the specific formulation provided by (5) may yield valuable insights applicable to a wide range of divergence measures. The following theorem provides lower and upper bounds for the UP function.
Theorem 3.
The uncertainty-perception function is confined to the following region
$\eta(P)\cdot N(X|Y)\,\leq\,U(P)\,\leq\,\eta(P)\cdot N(X_{G}|Y)$
where $1\leq\eta(P)\leq 2$ is a convex function w.r.t the perception index and is given by
$\eta(P)=\Big{(}2e^{\frac{2P}{d}}-\sqrt{(2e^{\frac{2P}{d}}-1)^{2}-1}\Big{)}.$
Noteworthy, Theorem 3 holds true regardless of the underlying distributions of $X$ and $Y$, thereby providing a universal characterization of the UP function in terms of perception. Furthermore, as depicted in Figure 2, Theorem 3 gives rise to the uncertainty-perception plane, which divides the space into three distinct regions:
  1. 1.
    Impossible region, where no estimator can reach.
  2. 2.
    Optimal region, encompassing all estimators that are optimal according to (5).
  3. 3.
    Suboptimal region of estimators which exhibit overly high uncertainty.
The existence of an impossible region highlights the uncertainty-perception tradeoff, proving no estimator can achieve both high perception and low uncertainty simultaneously. This finding underscores the importance of practitioners being aware of this tradeoff, enabling them to make informed decisions when prioritizing between perceptual quality and uncertainty in their applications. The uncertainty-perception plane could serve as a valuable framework for evaluating estimator performance in this context. While not a comprehensive metric, it may offer insights into areas where improvements can be made, guiding practitioners towards estimators that strike a more desirable balance between perception and uncertainty. For certain estimators residing in the suboptimal region, it may be possible to achieve lower uncertainty without sacrificing perceptual quality. Thus, we believe that our proposed uncertainty-perception plane can serve as a valuable starting point for further research and practical applications, ultimately leading to the development of safer and reliable image restoration algorithms.
Figure 2: The uncertainty-perception plane, derived from Theorem 3. The impossible region demonstrates the inherent tradeoff between perception and uncertainty, while other regions may guide practitioners toward estimators that balance perception and uncertainty, highlighting potential areas for improvement.
Next, we analyze how the dimensionality of the underlying data affects the uncertainty-perception tradeoff. To achieve this, we extend the function $\eta(P)$ to include a dimension parameter $d$, denoted as $\eta(P;d)$. As shown in Fig. 3, $\eta(P;d)$ exhibits a rapid incline as perception improves and it attain higher values in higher dimensions. This observation suggests that in high-dimensional settings, the uncertainty-perception tradeoff becomes more severe, implying that any marginal improvement in perception for an algorithm is accompanied by a dramatic increase in uncertainty.
Figure 3: Impact of dimensionality, as revealed in Theorem 3, demonstrates that the uncertainty-perception tradeoff intensifies in higher dimensions. This implies that even minor improvements in perceptual quality for an algorithm may come at the cost of a significant increase in uncertainty.
Finally, we conjecture that the general form of the tradeoff, given by the inequality in Theorem 3, holds for different divergence measures, with the specific form of $\eta(P)$ capturing the nuances of each chosen measure. For instance, considering the Hellinger distance as our perception measure, we obtain the same inequality as in Theorem 3 but with $\eta(P)$ defined for $0\leq P\leq 1$ as [^1] The case of $P=1$ is obtained by taking the limit $\underset{P\rightarrow 1}{\lim}\;\eta(P)=1$.
$\eta_{\text{Hellinger}}(P)=\frac{2}{(1-P)^{4/d}}-\sqrt{\left(\frac{2}{(1-P)^{4/d}}-1\right)^{2}-1}.$ (6)
4.2 Revisiting the Distortion-Perception Tradeoff
Having established the uncertainty-perception tradeoff and its characteristics, we now broaden our analysis to estimation distortion, particularly the mean squared-error. A well-known result in estimation theory states that for any random variable $X$ and for any estimator $\hat{X}$ based upon side information $Y$, the following holds true [19]:
$\mathbb{E}\left[||\hat{X}-X||^{2}\right]\geq\frac{1}{2\pi e}e^{2h(X|Y)}.$ (7)
This inequality, related to the uncertainty principle, serves as a fundamental limit to the minimal MSE achieved by any estimator. However, it does not consider the estimation uncertainty of $\hat{X}$ as the right hand side is independent of $\hat{X}$. Thus, we extend the above in the following theorem.
Theorem 4.
For any random variable $X$, observation $Y$ and unbiased estimator $\hat{X}$, it holds that
$\frac{1}{d}\mathbb{E}\left[||\hat{X}-X||^{2}\right]\geq N\left(\hat{X}-X\big{|}Y\right).$
Notice that for any estimator $\hat{X}$ we have $N(\hat{X}-X|Y)\geq N(X|Y)$, implying
$\frac{1}{d}\mathbb{E}[\|\hat{X}-X\|^{2}]\geq N(X|Y)=\frac{1}{2\pi e}e^{\frac{2}{d}h(X|Y)}.$ (8)
The above result aligns with equation (7), demonstrating that Theorem 4 serves as a generalization of inequality (7), incorporating the uncertainty associated with the estimation. Furthermore, by viewing the estimator $\hat{X}$ as a function of perception index $P$, we arrive at the next corollary.
Corollary 1.
Define the following distortion-perception function
$D(P)\triangleq\underset{p_{\hat{X}|Y}}{\min}\;\Big{\{}\frac{1}{d}\mathbb{E}\left[||\hat{X}-X||^{2}\right]:\;D_{v}(X,\hat{X}\big{|}Y)\leq P\Big{\}}.$
Then, for any perceptual index $P$, we have $D(P)\geq U(P)$.
As uncertainty increases with improving perception, the corollary implies that distortion also increases. Thus, when utilizing MSE as a measure of distortion, the uncertainty-perception tradeoff induces a distortion-perception tradeoff [12], offering a novel interpretation of this well-known phenomenon.
4.3 Quantitative Demonstration
Empirical observations are the catalyst for our theoretical framework. Thus, we support our theoretical findings via quantitative evaluation of image super-resolution (SR) algorithms, following previous works [12, 11, 24, 51]. Utilizing the BSD100 dataset [50], the goal is predicting a high-resolution (HR) image from its low-resolution counterpart obtained through 4×4$\times$4 × bicubic downsampling. We evaluate various recent SR algorithms, including EDSR [45], ESRGAN [71], SinGAN [63], SANGAN [38], DIP [67], SRResNet/SRGAN variants [43], EnhanceNet [58], and Latent Diffusion Models (LDMs) with parameter $\beta\in[0,1]$ [57], where $\beta=0$ recovers DDIM [33] and $\beta=1$ recovers DDPM [65]. Our performance analysis focuses on uncertainty, and common measures of perception and distortion. Perception measures include projected distribution loss (PDL) [22], LPIPS [75], and Single-FID (SFID) – a patch-based variation of Fréchet inception distance (FID) [32] for image pairs where we randomly crop patches from each image and compute the standard FID between the groups. Distortion measures include MSE, peak signal-to-noise ratio (PSNR), and structural similarity index (SSIM) [72]. Estimating entropy in high dimensions poses significant challenges [42]. Hence, we utilize an upper bound for uncertainty, $N(\hat{X}_{G}-X_{G}|Y)$, discussed in Appendix E. This practical alternative simplifies the computation to calculating the geometric mean of the singular values of the error covariance. Additional results using direct entropy estimation are presented in Appendix G.
Figure 4: Evaluation of SR algorithms. Top: Uncertainty-perception plane showing the tradeoff between perceptual quality and uncertainty (y-axis) for various perceptual measures. Bottom: Uncertainty-distortion plane illustrating the relation between uncertainty (x-axis) and distortion for different distortion measures. Axis placement differs in the two rows to highlight the distinct roles of uncertainty.
Figure 5: Visual comparison of algorithms on the uncertainty-perception plane. Algorithms are ordered from low to high uncertainty (left to right), with an accompanying increase in perceptual quality and distortion.
Figure 4 reveals the results of our analysis. As observed in the top row, across various perceptual measures, an unattainable blank region exists in the lower left corner, indicating that no model simultaneously achieves both low uncertainty and high perceptual quality. Furthermore, an anti-correlation emerges near this region, where modest improvements in perceptual quality translate to dramatic increases in uncertainty. This observation suggests the existence of a tradeoff between uncertainty and perception. Additionally, the bottom row showcases a strong relationship between uncertainty and distortion across diverse measures, demonstrating that any increase in uncertainty leads to a significant rise in distortion. [^2] Note that MSE is a measure of distortion, whereas PSNR and SSIM are measures of inverse distortion; this accounts for the negative slope in the first two figures, and the positive slope in the third. Figure 5 visually reinforces these findings by presenting selected recoveries ordered from low to high uncertainty. The results show a clear increase in perceptual quality with increasing uncertainty, accompanied by a visually subtle increase in distortion.
5 Conclusion
This study established the uncertainty-perception tradeoff in generative restoration, demonstrating that high perceptual quality leads to increased uncertainty, particularly in high dimensions. We characterized this tradeoff and its fundamental relation to the inherent uncertainty of the problem, introducing the uncertainty-perception plane which may guide practitioners in understanding estimator performance. By extending our analysis to MSE distortion, we showed that the distortion-perception tradeoff emerges as a direct consequence of the uncertainty-perception tradeoff. Experimental results confirmed our theoretical findings, highlighting the importance of this tradeoff in image restoration.
6 Limitations
While our theoretical analysis is applicable to any signal restoration task with non-invertible degradation, we focused our empirical validation on image super-resolution for consistency with prior art. Therefore, expanding the experiments to include additional image-to-image tasks and domains such as audio, video, and text could reveal broader implications and applications of this tradeoff.
Our analysis is grounded in the theoretical framework of entropy as a measure of uncertainty. However, the practical estimation of high-dimensional entropy remains an active area of research. To address this, we utilize a tractable upper bound in Section 4.3. This alternative approach calls for further study, investigating its potential to effectively quantify uncertainty in analyzing the performance of existing algorithms. Furthermore, incorporating this upper bound into the design of new algorithms, whether as an objective function to minimize or a regularization term to constrain, could enable explicit control over the uncertainty-perception tradeoff, potentially leading to more reliable solutions.
7 Broader Impact
Our work revealing a fundamental tradeoff between uncertainty and perception in image restoration carries significant societal impact. This knowledge empowers users and developers to utilize and build generative models more responsibly, particularly in critical fields like healthcare, forensics, and autonomous systems. By understanding this inherent trade-off, practitioners can consider trading performance for better safety and resilience against potential misuse and misinterpretations.
Appendix A Information-Theory Preliminaries
To make the paper self-contained, we briefly overview the essential definitions and results in information-theory. Let $X$, $Y$ and $Z$ be continuous random variables with probability density functions $p_{X}(x)$, $p_{Y}(y)$ and $p_{Z}(z)$ respectively. The space of probability density functions is denoted by $\Omega$. We assume the quantities described below, which involve integrals, are well-defined and finite.
Definition 3 (Entropy).
The differential entropy of $X$, whose support is a set $S_{x}$, is defined by
[…]
Definition 4 (Rényi Entropy).
The Rényi entropy of order $r\geq 0$ of $X$ is defined by
[…]
The above quantity generalizes various notions of entropy, including Hartley entropy, collision entropy, and min-entropy. In particular, for $r=1$ we have
$h_{1}(X)\triangleq\lim_{r\rightarrow 1}h_{r}(X)=h(X).$
Definition 5 (Entropy Power).
Let be $h(X)$ be the differential entropy of $X\in{\mathbb{R}}^{d}$. Then, the entropy Power of $X$ is given by
$N(X)\triangleq\frac{1}{2\pi e}e^{\frac{2}{d}h(X)}.$
Definition 6 (Divergence).
A statistical divergence is any function $D_{v}:\Omega\times\Omega\rightarrow{\mathbb{R}}^{+}$ which satisfies the following conditions for all $p,q\in\Omega$:
  1. 1.
    $D_{v}(p,q)\geq 0$.
  2. 2.
    $D_{v}(p,q)=0$ iff $p=q$ almost everywhere.
Table 1: Formulas for Multivariate Gaussian Distribution
Distribution Quantity Closed-Form Expression
$X\sim{\mathcal{N}}(\mu_{x},\Sigma_{x})$ $h(X)$ $\frac{1}{2}\ln\{(2\pi e)^{d}\left|\Sigma_{x}\right|\}.$
$X\sim{\mathcal{N}}(\mu_{x},\Sigma_{x})$ $N(X)$ $\left|\Sigma_{x}\right|^{1/n}.$
$X\sim{\mathcal{N}}(\mu_{x},\Sigma_{x})$ $h_{\frac{1}{2}}(X)$ $\frac{1}{2}\ln\{(8\pi)^{d}\left|\Sigma_{x}\right|\}.$
$X\sim{\mathcal{N}}(\mu_{x},\Sigma_{x})$, $Y\sim{\mathcal{N}}(\mu_{y},\Sigma_{y})$ $D_{1/2}(X,Y)$ $\frac{1}{4}(\mu_{x}-\mu_{y})^{T}\left(\frac{\Sigma_{x}+\Sigma_{y}}{2}\right)^{-1}(\mu_{x}-\mu_{y})+\ln\left(\frac{\left|\frac{\Sigma_{x}+\Sigma_{y}}{2}\right|}{\sqrt{\left|\Sigma_{x}\right|\left|\Sigma_{y}\right|}}\right).$
Definition 7 (Rényi Divergence).
The Rényi divergence of order $r\geq 0$ between $p_{X}$ and $p_{Y}$ is
[…]
The above establishes a spectrum of divergence measures, generalising the Kullback–Leibler divergence as $D_{1}(X,Y)=D_{KL}(X,Y)$. Furthermore, it is important to note that all orders $r\in(0,1)$ are equivalent [69], since
[…]
Definition 8 (Conditioning).
Consider the joint probability $p_{XY}$ and the conditional probabilities $p_{X|Y}(x|y)$ and $p_{Z|Y}(z|y)$. The conditional differential entropy of $X\in{\mathbb{R}}^{d}$ given $Y$ is defined as
[…]
$=\mathbb{E}_{y\sim p_{Y}}\left[h(X|Y=y)\right]$
where $S_{XY}$ is the support set of $p_{XY}$. Then, the conditional entropy power of $X$ given $Y$ is
$N(X|Y)=\frac{1}{2\pi e}e^{\frac{2}{d}h(X|Y)}.$
Similarly, the conditional divergence between $X$ and $Z$ given $Y$ is defined as
$D_{v}(X,Z\big{|}Y)\triangleq\mathbb{E}_{y\sim p_{Y}}\left[D_{v}(X|Y=y,Z|Y=y)\right].$
For example, the conditional Rényi divergence is given by
$D_{r}(X,$ $Z\big{|}Y)\triangleq$
[…]
Table 1 summarizes closed-form expressions for several quantities relevant to the multivariate Gaussian distribution. Below we present two fundamental results that form the basis of our analysis.
Lemma 1 (Maximum Entropy Principle [19]).
Let $X\in{\mathbb{R}}^{d}$ be a continuous random variable with zero mean and covariance $\Sigma_{x}$. Define $X_{G}\sim\mathcal{N}(0,\Sigma_{x})$ to be a Gaussian random variable, independent of $X$, with the identical covariance matrix $\Sigma_{x_{G}}=\Sigma_{x}$. Then,
$h(X)\leq h(X_{G}),$
$N(X)\leq N(X_{G})=\left|\Sigma_{x}\right|^{1/d}.$
Lemma 2 (Entropy Power Inequality [48]).
Let $X$ and $Y$ be independent continuous random variables. Then, the following inequality holds
$N(X)+N(Y)\leq N(X+Y),$
where equality holds iff $X$ and $Y$ are multivariate Gaussian random variables with proportional covariance matrices. Equivalently, let $X_{g}$ and $Y_{g}$ be defined as independent, isotropic multivariate Gaussian random variables satisfying $h(X_{g})=h(X)$ and $h(Y_{g})=h(Y)$. Then,
$h(X)+h(Y)=h(X_{g})+h(Y_{g})=h(X_{g}+Y_{g})\leq h(X+Y).$
Appendix B Derivation of Example 1
Since $\hat{X}=\mathbb{E}\left[X|Y\right]+Z$, then $\hat{X}|Y\sim{\mathcal{N}}(\mathbb{E}\left[X|Y\right],\sigma_{z}^{2})$. Moreover, $X|Y\sim{\mathcal{N}}(\mathbb{E}\left[X|Y\right],\sigma_{q}^{2})$ where $\sigma_{q}^{2}=\frac{\sigma 2}{1+\sigma^{2}}$. Thus, the conditional error entropy is given by $N(\hat{X}-X|Y)=\sigma_{q}^{2}+\sigma_{z}^{2}$ and the symmetric KL divergence is $D_{SKL}(X,\hat{X}\big{|}Y)=\frac{\sigma_{q}^{2}+\sigma_{z}^{2}}{2\sigma_{z}\sigma_{q}}-1$, leading the following problem
$U(P)=\underset{\sigma_{z}}{\min}\;\Big{\{}\sigma_{q}^{2}+\sigma_{z}^{2}\;:\;\frac{\sigma_{q}^{2}+\sigma_{z}^{2}}{2\sigma_{z}\sigma_{q}}-1\leq P\Big{\}}.$ (10)
Therefore, we seek the minimal value of $\sigma_{z}$ that satisfies the constraint. Note that the minimal value is attained at the boundary of the constraint set, where the inequality becomes an equality
$\begin{split}\frac{\sigma_{q}^{2}+\sigma_{z}^{2}}{2\sigma_{z}\sigma_{q}}-1=P\;\Rightarrow\;\sigma_{z}^{2}-2\sigma_{q}(P+1)\sigma_{z}+\sigma_{q}^{2}=0.\end{split}$ (11)
The solution to the aforementioned quadratic problem is $\sigma_{z}^{*}=\sigma_{q}\left(P+1-\sqrt{(P+1)^{2}-1}\right)$. Substituting the later into the objective function, we obtain
$U(P)=\sigma_{q}^{2}\Big{[}1+\left(P+1-\sqrt{(P+1)^{2}-1}\right)^{2}\Big{]}.$ (12)
Finally, the entropy power of an univariate Gaussian distribution equals its variance $\sigma_{q}^{2}=N(X|Y)$. Figure 6 visualizes the resulting uncertainty-perception tradeoff.
Figure 6: The Uncertainty-Perception function for Example 1. As perception quality improves, the minimal achievable uncertainty increases, suggesting a tradeoff governed by the inherent uncertainty.
Appendix C Proof of Theorem 1
First, the constraint ${\mathcal{C}}(P)\triangleq\{\hat{X}\,:\,D_{v}(X,\hat{X}\big{|}Y)\leq P$} defines a compact set which is continuous in $P$. Hence, by the Maximum Theorem [19], $U(P)$ is continuous. In addition, $U(P)$ is the minimal error entropy power obtained over a constraint set whose size does not decrease with $P$, thus, $U(P)$ is non-increasing in $P$. Any continuous non-increasing function is quasi-linear. For the lower bound consider the case where $P=\infty$, leading to the following unconstrained problem
$U(\infty)\triangleq\underset{p_{\hat{X}|Y}}{\min}\;N(\hat{X}-X|Y).$ (13)
For any $P\geq 0$ it holds that $U(\infty)\leq U(P)$, and by Lemma 2 we have
$N(X|Y)+\min_{p_{\hat{X}|Y}}\;N(\hat{X}|Y)\leq U(\infty).$ (14)
Since $\min_{p_{\hat{X}|Y}}\;N(\hat{X}|Y)\geq 0$ we obtain
$\forall P\geq 0:\quad N(X|Y)\leq U(P).$ (15)
Next, we have $U(P)\leq U(0)=N(\hat{X}_{0}-X|Y)$ where $p_{\hat{X}_{0}|Y}=p_{X|Y}$. Define $V\triangleq\hat{X}_{0}-X$, then $\Sigma_{v|y}=\Sigma_{\hat{x}|y}+\Sigma_{x|y}=2\Sigma_{x|y}$ where we use that $X$ and $\hat{X}$ are independent given $Y$. Thus,
$U(0)=N(V|Y)\leq N(V_{G}|Y)=\left|\Sigma_{v|y}\right|^{1/d}=\left|2\Sigma_{x|y}\right|^{1/d}=2\left|\Sigma_{x|y}\right|^{1/d}=2N(X_{G}|Y),$ (16)
where the first inequality is due to Lemma 1. Finally, for any $P\geq 0$ it holds that $U(P)\leq U(0)$ which implies $U(0)\leq 2N(X_{G}|Y)$, completing the proof.
Appendix D Proof of Theorem 2
Assuming $D_{v}(X,\hat{X}\big{|}Y)$ is convex in its second argument, the constraint represent a compact, convex set. Moreover, $h(\hat{X}-X|Y)$ is strictly-concave w.r.t $p_{\hat{X}|Y}$ as a composition of a linear function (convolution) with a strictly-concave function (entropy). Therefore, we minimize a log-concave function over a convex domain and thus the global minimum is attained on the set boundary where $D_{v}(X,\hat{X}\big{|}Y)=P$.
Appendix E Proof of Theorem 3
We begin with applying Lemma 1 and Lemma 2 to bound the objective function as follows
$N(\hat{X}_{g}|Y)+N(X_{g}|Y)=N(\hat{X}_{g}-X_{g}|Y)\leq N(\hat{X}-X|Y)\leq N(\hat{X}_{G}-X_{G}|Y).$ (17)
Note that the bounds are tight as the upper bound is attained when $\hat{X}|Y$ and $X|Y$ are multivariate Gaussian random variables, while the lower bound is attained if we further assume they are isotropic. Thus, we can bound the uncertainty-perception function as follows
$U_{g}(P)\leq U(P)\leq U_{G}(P)$ (18)
where we define
$\begin{split}&U_{g}(P)\triangleq\underset{p_{\hat{X}_{g}|Y}}{\min}\;\Big{\{}N(\hat{X}_{g}|Y)+N(X_{g}|Y)\;:\;D_{1/2}(X_{g},\hat{X}_{g}\big{|}Y)\leq P\Big{\}},\\ &U_{G}(P)\triangleq\underset{p_{\hat{X}_{G}|Y}}{\min}\;\Big{\{}N(\hat{X}_{G}-X_{G}|Y)\;:\;D_{1/2}(X_{G},\hat{X}_{G}\big{|}Y)\leq P\Big{\}}.\end{split}$ (19)
The above quantities can be expressed in closed form. We start with minimization problem of the upper bound which can be written as
$U_{G}(P)=\underset{p_{\hat{X}_{G}|Y}}{\min}\;\Big{\{}\frac{1}{2\pi e}e^{\frac{2}{d}\mathbb{E}\left[h(\hat{X}_{G}-X_{G}|Y=y)\right]}\;:\;\mathbb{E}\left[D_{1/2}(X_{G},\hat{X}_{G}\big{|}Y=y)\right]\leq P\Big{\}},$ (20)
where the expectation is over $y\sim Y$. Substituting the expressions for $h(X_{G}-X_{G}|Y=y)$ and $D_{1/2}(X_{G},\hat{X}_{G}\big{|}Y=y)$, we get
$U_{G}(P)=\underset{\{\Sigma_{\hat{x}|y}\}}{\min}\;\Bigg{\{}\frac{1}{2\pi e}e^{\frac{2}{d}\mathbb{E}\left[\frac{1}{2}\log\Big{\{}(2\pi e)^{d}\left|\Sigma_{\hat{x}|y}+\Sigma_{x|y}\right|\Big{\}}\right]}\;:\;\mathbb{E}\left[\log\frac{\left|\left(\Sigma_{\hat{x}|y}+\Sigma_{x|y}\right)/2\right|}{\sqrt{\left|\Sigma_{\hat{x}|y}\right|\left|\Sigma_{x|y}\right|}}\right]\leq P\Bigg{\}}.$ (21)
Notice the optimization is with respect to the covariance matrices $\{\Sigma_{\hat{x}|y}\}$. Simplifying the above, we can equivalently solve the following minimization
$\underset{\{\Sigma_{\hat{x}|y}\}}{\min}\;\mathbb{E}\left[\log\left|\Sigma_{\hat{x}|y}+\Sigma_{x|y}\right|\right]\;\text{s.t.}\;\mathbb{E}\left[\log\frac{\left|\left(\Sigma_{\hat{x}|y}+\Sigma_{x|y}\right)/2\right|}{\sqrt{\left|\Sigma_{\hat{x}|y}\right|\left|\Sigma_{x|y}\right|}}\right]\leq P.$ (22)
The solution of a constrained optimization problem can be found by minimization the Lagrangian
$L\left(\{\Sigma_{\hat{x}|y}\},\lambda\right)\triangleq\mathbb{E}\left[\log\left|\Sigma_{\hat{x}|y}+\Sigma_{x|y}\right|\right]+\lambda\left(\mathbb{E}\left[\log\frac{\left|\left(\Sigma_{\hat{x}|y}+\Sigma_{x|y}\right)/2\right|}{\sqrt{\left|\Sigma_{\hat{x}|y}\right|\left|\Sigma_{x|y}\right|}}\right]-P\right).$ (23)
Since expectation is a linear operation and using that $P=\mathbb{E}\left[P\right]$, we rewrite the above as
$L\left(\{\Sigma_{\hat{x}|y}\},\lambda\right)=\mathbb{E}\left[\log\left|\Sigma_{\hat{x}|y}+\Sigma_{x|y}\right|+\lambda\left(\log\frac{\left|\left(\Sigma_{\hat{x}|y}+\Sigma_{x|y}\right)/2\right|}{\sqrt{\left|\Sigma_{\hat{x}|y}\right|\left|\Sigma_{x|y}\right|}}-P\right)\right].$ (24)
The expression within the expectation can be written as
$\log\left|\Sigma_{\hat{x}|y}+\Sigma_{x|y}\right|+\lambda\left(\log\left|\left(\Sigma_{\hat{x}|y}+\Sigma_{x|y}\right)/2\right|-\frac{1}{2}\log\left|\Sigma_{\hat{x}|y}\right|-\frac{1}{2}\log\left|\Sigma_{x|y}\right|-P\right).$ (25)
Next, according to KKT conditions the solutions should satisfy $\frac{\partial L}{\partial\Sigma_{\hat{x}|y}}=0$. Using the linearity of the expectation and differentiating (25) w.r.t $\Sigma_{\hat{x}|y}$ we obtain
$\left(\Sigma_{\hat{x}|y}+\Sigma_{x|y}\right)^{-1}+\lambda\left(\left(\Sigma_{\hat{x}|y}+\Sigma_{x|y}\right)^{-1}-\frac{1}{2}\Sigma_{\hat{x}|y}^{-1}\right)=0$ (26)
Multiplying both sides by $\left(\Sigma_{\hat{x}|y}+\Sigma_{x|y}\right)$, we have
$\begin{split}&I+\lambda I-\frac{\lambda}{2}I-\frac{\lambda}{2}\Sigma_{x|y}\Sigma_{\hat{x}|y}^{-1}=0\\ \Rightarrow\;&(1+\frac{\lambda}{2})I=\frac{\lambda}{2}\Sigma_{x|y}\Sigma_{\hat{x}|y}^{-1}\\ \Rightarrow\;&(\lambda+2)\Sigma_{\hat{x}|y}=\lambda\Sigma_{x|y}\\ \Rightarrow\;&\Sigma_{\hat{x}|y}=\frac{\lambda}{\lambda+2}\Sigma_{x|y}.\end{split}$ (27)
Define $\gamma=\frac{\lambda}{\lambda+2}$, so $\Sigma_{\hat{x}|y}=\gamma\Sigma_{x|y}$. Substituting the latter into the constraint we get
$\begin{split}&\log\left|\left(\gamma\Sigma_{x|y}+\Sigma_{x|y}\right)/2\right|-\frac{1}{2}\log\left|\gamma\Sigma_{x|y}\right|-\frac{1}{2}\log\left|\Sigma_{x|y}\right|=P\\ \Rightarrow\;&n\log\frac{1+\gamma}{2}-\frac{n}{2}\log\gamma=P\\ \Rightarrow\;&\frac{(1+\gamma)^{2}}{4\gamma}=e^{\frac{2}{d}P}\\ \Rightarrow\;&\gamma^{2}+2\gamma+1=4\gamma e^{\frac{2}{d}P}\\ \Rightarrow\;&\gamma(P)=2e^{\frac{2}{d}P}-1-\sqrt{(2e^{\frac{2}{d}P}-1)^{2}-1}.\end{split}$ (28)
Thus, we obtain that
$U_{G}(P)=\eta(P)\cdot N(X_{G}|Y)$ (29)
where
$\eta(P)=\gamma(P)+1=2e^{\frac{2}{d}P}-\sqrt{(2e^{\frac{2}{d}P}-1)^{2}-1}.$ (30)
Notice that $\eta(0)=2$, while $\lim_{P\rightarrow\infty}\;\eta(P)=1$, so $1\leq\eta(P)\leq 2$. Following similar steps where we replace $\Sigma_{\hat{x}|y}$ and $\Sigma_{x|y}$ with $N(\hat{X}|Y)$ and $N(X|Y)$ respectively, we derive
$U_{g}(P)=\eta(P)\cdot N(X|Y).$ (31)
Appendix F Proof of Theorem 4
Define ${\mathcal{E}}\triangleq\hat{X}-X$. Then,
$\frac{1}{d}\mathbb{E}\left[||\hat{X}-X||^{2}\right]$ =$\underset{(a)}{=}\mathbb{E}\left[\frac{1}{d}\mathbb{E}\left[||\hat{X}-X||^{2}\big{|}Y\right]\right]=\mathbb{E}\left[\frac{1}{d}\mathbb{E}\left[||{\mathcal{E}}||^{2}\big{|}Y\right]\right]=\mathbb{E}\left[\frac{1}{d}\mathbb{E}\left[{\mathcal{E}}^{T}{\mathcal{E}}\big{|}Y\right]\right]$
$=\mathbb{E}\left[\frac{1}{d}Tr\left(\mathbb{E}\left[{\mathcal{E}}{\mathcal{E}}^{T}\big{|}Y\right]\right)\right]=\mathbb{E}\left[\frac{1}{d}Tr\left(\Sigma_{\varepsilon|y}\right)\right]$
$\underset{(b)}{\geq}\mathbb{E}\left[\left|\Sigma_{\varepsilon|y}\right|^{1/d}\right]=\mathbb{E}\left[\left|\Sigma_{\hat{x}|y}+\Sigma_{x|y}\right|^{1/d}\right]$
$\underset{(c)}{\geq}\mathbb{E}\left[\frac{1}{2\pi e}e^{\frac{2}{d}h(\hat{X}-X|Y=y)}\right]$
$\underset{(d)}{\geq}\frac{1}{2\pi e}e^{\frac{2}{d}\mathbb{E}\left[h(\hat{X}-X|Y=y)\right]}=\frac{1}{2\pi e}e^{\frac{2}{d}h(\hat{X}-X|Y)}=N\left(\hat{X}-X\big{|}Y\right),$
where (a) is by the law of total expectation, (b) is due to the inequality of arithmetic and geometric means, (c) follows Lemma 1, and (d) is according to Jensen’s inequality.
Appendix G Results via Direct Estimation
Estimating high-dimensional statistics is prone to errors [42]. we used practical measures for perceptual quality and a tractable upper bound for uncertainty. Here, we supplement those results with direct computations of entropy and divergence in a high-dimensional setting. Following prior work [12, 24], we treat images as stationary sources and extract $9\times 9$ patches. To estimate Rényi divergence for perceptual quality assessment, we first model the probability density functions using kernel density estimation. Subsequently, we compute the divergence through empirical expectations. Uncertainty is estimated using the Kozachenko-Leonenko estimator, which calculates the patch sample differential entropy based on nearest neighbor distances [39, 21, 9, 49]. Results, shown in Figure 7, strongly align with the trends observed in Figure 4.
Figure 7: Evaluation of SR algorithms via direct estimation of high-dimensional statistics. Left: Uncertainty-perception plane demonstrating the tradeoff between perceptual quality and uncertainty. Right: Uncertainty-distortion plane illustrating the relation between uncertainty and distortion. Results are consistent with the finding in Figure 4.

### v3

@@
 However, as their perceptual quality continues to improve, these models also exhibit a growing tendency to generate hallucinations – realistic-looking details that do not exist in the ground truth images.
-The presence of hallucinations introduces uncertainty regarding the reliability of the models’ predictions, raising major concerns about their practical application.
-In this paper, we employ information-theory tools to investigate this phenomenon, revealing a fundamental tradeoff between uncertainty and perception.
+Hallucinations in these models create uncertainty about their reliability, raising major concerns about their practical application.
+This paper investigates this phenomenon through the lens of information theory, revealing a fundamental tradeoff between uncertainty and perception.
 We rigorously analyze the relationship between these two factors, proving that the global minimal uncertainty in generative models grows in tandem with perception.
 In particular, we define the inherent uncertainty of the restoration problem and show that attaining perfect perceptual quality entails at least twice this uncertainty.
-Additionally, we establish a relation between mean squared-error distortion, uncertainty and perception, through which we prove the aforementioned uncertainly-perception tradeoff induces the well-known perception-distortion tradeoff.
+Additionally, we establish a relation between distortion, uncertainty and perception, through which we prove the aforementioned uncertainly-perception tradeoff induces the well-known perception-distortion tradeoff.
+We demonstrate our theoretical findings through experiments with super-resolution and inpainting algorithms.
 This work uncovers fundamental limitations of generative models in achieving both high perceptual quality and reliable predictions for image restoration.
-We demonstrate our theoretical findings through an analysis of single image super-resolution algorithms.
-Our work aims to raise awareness among practitioners about this inherent tradeoff, empowering them to make informed decisions and potentially prioritize safety over perceptual performance.
+Thus, we aim to raise awareness among practitioners about this inherent tradeoff, empowering them to make informed decisions and potentially prioritize safety over perceptual performance.
+Figure 1: Illustration of Theorem 3.
+In restoration tasks, the minimal attainable uncertainty is lower bounded by a function that begins at the inherent uncertainty $U_{\text{Inherent}}$ of the problem (Definition 2) and graudally increases up to twice this value as the recovery approaches perfect perceptual quality.
 1 Introduction
 Restoration tasks and inverse problems impact many scientific and engineering disciplines, as well as healthcare, education, communication and art.
-Generative artificial intelligence has transformed the field of inverse problems due to its unprecedented ability to infer missing information and restore corrupted data.
-In the realm of image restoration, the quest for high perceptual quality has led to a new generation of generative models  [33, 65, 18, 17, 70], capable of producing outputs of remarkable realism, virtually indistinguishable from true images.
+Generative artificial intelligence [80, 38, 10] has transformed the field of inverse problems due to its unprecedented ability to infer missing information and restore corrupted data.
+In the realm of image restoration, the quest for high perceptual quality has led to a new generation of generative models, capable of producing outputs of remarkable realism, virtually indistinguishable from true images.
 While powerful, growing empirical evidence indicates that generative models are susceptible to hallucinations [30], characterized by the generation of seemingly authentic content that deviates from the original input data, hindering applications where faithfulness is crucial.
@@
 These approaches offer crucial insights into the model’s confidence in its predictions, empowering users to assess potential deviations from the original data and make informed decisions.
-Despite these advancements, the intricate relationship between achieving high perceptual quality and the extent of uncertainty remains an understudied area.
+Despite this progress, the relationship between achieving high perceptual quality and the extent of uncertainty remains an understudied area.
 This paper establishes the theoretical relationship between uncertainty and perception, demonstrating through rigorous analysis that the global minimal uncertainty in generative models increases with the level of desired perceptual quality (see illustration in Figure 1).
@@
   3. 3.
-    We establish a relationship between uncertainty and mean squared error (MSE) distortion, demonstrating that the uncertainty-perception trade-off induces the well-known distortion-perception trade-off (Theorem 4).
-We empirically validate all theoretical findings through experiments with image super-resolution algorithms (Section 4.3).
+    We establish a relationship between uncertainty and mean squared error (MSE) distortion, demonstrating that the uncertainty-perception trade-off induces the well-known distortion-perception trade-off [14] (Theorem 4).
+  4. 4.
+    We empirically validate all theoretical findings through experiments on image super-resolution and inpainting (Section 5), covering a broad spectrum of recovery algorithms, diverse metrics and data distributions.
+Our experimental results for image inpainting are illustrated in Figure 2.
+Figure 2: Image inpainting results. Algorithms are ordered from low to high perception (left to right). Note the corresponding increased hallucinations and distortion. See Section 5 for details.
 We aim to provide practitioners with a deeper understanding of the tradeoff between uncertainty and perceptual quality, allowing them to strategically navigate this balance and prioritize safety when deploying generative models in real-world, sensitive applications.
-Figure 1: Illustration of Theorem 3.
-In inverse problems, the minimal attainable uncertainty exhibits a lower bound that begins at the inherent uncertainty $U_{\text{Inherent}}$ of the problem (Definition 2) and increases up to twice this value as we approach perfect perceptual quality of the recovery.
 2 Related Work
@@
 Recently, researchers have leveraged image-level embeddings from large vision models like DINO [17] and CLIP [62] to capture high-level similarity.
+Further advancements include HyperIQA [74] that leverages self-adaptive hyper networks to blindly assess image quality in the wild, while LIQE [88] and QAlign [84] utilize large language models to capture high-level semantic similarity and alignment between the restored and original images.
 Here, we follow previous works [58, 14, 31] and adopt a mathematical notion of perceptual quality defined as the divergence between probability densities.
@@
 In [58], it is shown that the Lipschitz constant of any deterministic estimator grows to infinity as it approaches perfect perception.
-Our work centers on uncertainty, reflecting the confidence level of a recovery algorithm.
-Thus, our works complements these studies by introducing uncertainty into the equation and exploring its impact on the perception-distortion relationship.
+This work uniquely emphasizes uncertainty in image restoration, distinguishing it from distortion.
+While distortion measures how close a restored image is to the original, uncertainty quantifies the confidence in the restoration itself.
+This distinction is crucial for decision-making, as high uncertainty can hinder informed choices, complementing existing research on perceptual quality and robustness.
 3 Problem Formulation
@@
 In this paper, we analyze estimators $\hat{X}$ with respect to two performance criteria: perception and uncertainty.
-To assess perceptual quality, we follow a theoretical approach, similar to previous works [74, 12], and measure perception using conditional divergence between $X$ and $\hat{X}$ defined as
+To assess perceptual quality, we follow a theoretical approach, similar to previous works [85, 14], and measure perception using conditional divergence [^1] See Appendix A for a brief explanation of how conditional divergence relates to human perception.
+between $X$ and $\hat{X}$ defined as
 $D_{v}(X,\hat{X}\big{|}Y)\triangleq\mathbb{E}_{y\sim p_{Y}}\left[D_{v}\big{(}p_{X|Y=y},p_{\hat{X}|Y=y}\big{)}\right],$ (1)
@@
 Therefore, we adopt the concept of entropy power from information theory, which assesses the statistical spread of a random variable.
-For the definition of entropy power and other relevant background, we refer the reader to Appendix A.
+For the definition of entropy power and other relevant background, we refer the reader to Appendix B.
 Utilizing entropy power, we formally define the inherent uncertainty intrinsic to the restoration problem as follows
@@
 Let the perception measure be the symmetric Kullback–Leibler (KL) divergence $D_{\text{SKL}}$ and assume stochastic estimators of the form $\hat{X}=\mathbb{E}\left[X|Y\right]+Z$ where $Z\sim{\mathcal{N}}(0,\sigma_{z}^{2})$ is independent of $Y$.
-As derived in Appendix B, the UP function admits a closed form expression in this case, given by
+As derived in Appendix C, the UP function admits a closed form expression in this case, given by
 $U(P)=N(X|Y)\Big{[}1+\left(P+1-\sqrt{(P+1)^{2}-1}\right)^{2}\Big{]},\text{ where }N(X|Y)=\sigma^{2}/(1+\sigma^{2}).$
-The above result, illustrated in Appendix B, demonstrates the minimal attainable uncertainty increases as the perception quality improves.
+The above result, illustrated in Appendix C, demonstrates the minimal attainable uncertainty increases as the perception quality improves.
 Moreover, The above example suggests a structure for uncertainty-perception function $U(P)$, which fundamentally relies on the inherent uncertainty $N(X|Y)$.
@@
 This tradeoff is fundamentally linked to the inherent uncertainty $N(X|Y)$ arising from the information loss during the observation process.
-Notably, the upper bound of $N(X|Y)$ can be expressed as
+Notably, the upper bound can be expressed as
 $N(X_{G}|Y)=N(X|Y)e^{\frac{2}{d}D_{KL}(X,X_{G}|Y)}.$ (4)
 This shows that as $X$ approaches Gaussianity, $N(X|Y)$ approaches $N(X_{G}|Y)$.
-However, concurrently, it implies in general higher values of $N(X|Y)$ due to Lemma 1 of Appendix A.
+However, concurrently, it implies in general higher values of $N(X|Y)$ due to Lemma 1 of Appendix B.
 This finding yields a surprising insight: for multivariate Gaussian distributions, perfect perceptual quality comes at the expense of exactly twice the inherent uncertainty of the problem.
-The next theorem shows that for a fixed perceptual index $P$, the optimal algorithms lie on the boundary of the constraint set.
-This facilitates the optimization task, as it restricts the search space to the boundary points.
+Next, we show that for a fixed perceptual index $P$, the optimal algorithms lie on the boundary of the constraint set.
+This facilitates the optimization, as it restricts the search space to the boundary points.
 Theorem 2.
 Assume $D_{v}(X,\hat{X}\big{|}Y)$ is convex in its second argument. Then, for any $P\geq 0$, the minimum is attained on the boundary where $D_{v}(X,\hat{X}\big{|}Y)=P$.
+Note that the assumption of the convexity of $\mathcal{D}_{v}$ in its second argument is not a restrictive condition.
+In fact, most widely-used divergence functions, notably all $f$-divergences (such as KL divergence, total variation distance, Hellinger distance, and Chi-square divergence), exhibit this property.
 While the above theorems describe important characteristics of the uncertainty-perception function, additional assumptions are needed to gain deeper insights.
@@
 $U(P)=\underset{p_{\hat{X}|Y}}{\min}\Big{\{}N(\hat{X}-X|Y)\;:\;D_{1/2}(X,\hat{X}\big{|}Y)\leq P\Big{\}}.$ (5)
-While we set $r=1/2$ to facilitate our derivations, it is important to note that all orders $r\in(0,1)$ are equivalent (see Appendix A).
+While we set $r=1/2$ to facilitate our derivations, it is important to note that all orders $r\in(0,1)$ are equivalent (see Appendix B).
 Consequently, given this equivalence and the close relationship between Rényi divergence and other metrics, analyzing the specific formulation provided by (5) may yield valuable insights applicable to a wide range of divergence measures.
@@
 Thus, we believe that our proposed uncertainty-perception plane can serve as a valuable starting point for further research and practical applications, ultimately leading to the development of safer and reliable image restoration algorithms.
-Figure 2: The uncertainty-perception plane, derived from Theorem 3.
-The impossible region demonstrates the inherent tradeoff between perception and uncertainty, while other regions may guide practitioners toward estimators that balance perception and uncertainty, highlighting potential areas for improvement.
+Figure 3: The uncertainty-perception plane (Theorem 3).
+The impossible region demonstrates the inherent tradeoff between perception and uncertainty, while other regions may guide practitioners toward estimators that better balance the two factors, highlighting potential areas for improvement.
 Next, we analyze how the dimensionality of the underlying data affects the uncertainty-perception tradeoff.
@@
 As shown in Fig.
-3, $\eta(P;d)$ exhibits a rapid incline as perception improves and it attain higher values in higher dimensions.
+4, $\eta(P;d)$ exhibits a rapid incline as perception improves and it attain higher values in higher dimensions.
 This observation suggests that in high-dimensional settings, the uncertainty-perception tradeoff becomes more severe, implying that any marginal improvement in perception for an algorithm is accompanied by a dramatic increase in uncertainty.
@@
 Thus, when utilizing MSE as a measure of distortion, the uncertainty-perception tradeoff induces a distortion-perception tradeoff [14], offering a novel interpretation of this well-known phenomenon.
-4.3 Quantitative Demonstration
-Empirical observations are the catalyst for our theoretical framework.
-Thus, we support our theoretical findings via quantitative evaluation of image super-resolution (SR) algorithms, following previous works [12, 11, 24, 51].
-Utilizing the BSD100 dataset [50], the goal is predicting a high-resolution (HR) image from its low-resolution counterpart obtained through 4×4$\times$4 × bicubic downsampling.
-We evaluate various recent SR algorithms, including EDSR [45], ESRGAN [71], SinGAN [63], SANGAN [38], DIP [67], SRResNet/SRGAN variants [43], EnhanceNet [58], and Latent Diffusion Models (LDMs) with parameter $\beta\in[0,1]$ [57], where $\beta=0$ recovers DDIM [33] and $\beta=1$ recovers DDPM [65].
-Our performance analysis focuses on uncertainty, and common measures of perception and distortion.
-Perception measures include projected distribution loss (PDL) [22], LPIPS [75], and Single-FID (SFID) – a patch-based variation of Fréchet inception distance (FID) [32] for image pairs where we randomly crop patches from each image and compute the standard FID between the groups.
-Distortion measures include MSE, peak signal-to-noise ratio (PSNR), and structural similarity index (SSIM) [72].
-Estimating entropy in high dimensions poses significant challenges [42].
-Hence, we utilize an upper bound for uncertainty, $N(\hat{X}_{G}-X_{G}|Y)$, discussed in Appendix E.
-This practical alternative simplifies the computation to calculating the geometric mean of the singular values of the error covariance.
-Additional results using direct entropy estimation are presented in Appendix G.
+5 Experiments
+Setup.
+Our theoretical framework is grounded in empirical observations, leading us to validate our findings through experiments on common benchmark tasks: image super-resolution and inpainting.
+We analyze performance through the lens of uncertainty, alongside established measures of perceptual quality and distortion.
+To assess perceptual quality, we employ state-of-the-art metrics including HyperIQA [74], LIQE [88] and Q-ALIGN [84].
+Distortion is evaluated using traditional measures: MSE, peak signal-to-noise ratio (PSNR), and structural similarity index (SSIM) [82].
+Accurately estimating entropy in high-dimensional spaces presents significant challenges [46]; hence, we utilize an upper bound for uncertainty, $N(\hat{X}_{G}-X_{G}|Y)$, as detailed in Appendix F.
+This practical alternative simplifies computation to calculating the geometric mean of the singular values of the error covariance.
+For super-resolution, we utilize the BSD100 benchmark dataset [55], aiming to predict a high-resolution image from its low-resolution counterpart obtained via 4×4$\times$4 × bicubic downsampling.
+Our evaluation spans a diverse range of recovery algorithms, including EDSR [50], ESRGAN [81], SinGAN [71], SANGAN [39], DIP [77], SRResNet/SRGAN variants [47], EnhanceNet [66], and Latent Diffusion Models (LDMs) with parameter $\beta\in[0,1]$ [65], where $\beta=0$ recovers DDIM [32] and $\beta=1$ recovers DDPM [73].
+In the context of image inpainting, we leverage the SeeTrue dataset [86], an image-text alignment benchmark known for its diverse collection of real and synthetic text-image pairs.
+Here, we focus our analysis on diffusion models due to their state-of-the-art performance and growing popularity in the field.
 Figure 5: Evaluation of SR algorithms.
 Top: Uncertainty-perception plane showing the tradeoff between perceptual quality and uncertainty (y-axis) for various perceptual measures.
-Bottom: Uncertainty-distortion plane illustrating the relation between uncertainty (x-axis) and distortion for different distortion measures.
+Bottom: Uncertainty-distortion plane showing the relationship between uncertainty and various distortion measures.
 Axis placement differs in the two rows to highlight the distinct roles of uncertainty.
-Figure 5: Visual comparison of algorithms on the uncertainty-perception plane.
-Algorithms are ordered from low to high uncertainty (left to right), with an accompanying increase in perceptual quality and distortion.
-Figure 4 reveals the results of our analysis.
-As observed in the top row, across various perceptual measures, an unattainable blank region exists in the lower left corner, indicating that no model simultaneously achieves both low uncertainty and high perceptual quality.
+Results.
+Figure 5 presents our super-resolution analysis.
+As observed in the top row, across various perceptual measures, an unattainable blank region exists in the lower right corner, indicating that no model simultaneously achieves both low uncertainty and high perceptual quality.
 Furthermore, an anti-correlation emerges near this region, where modest improvements in perceptual quality translate to dramatic increases in uncertainty.
@@
 [^3] Note that MSE is a measure of distortion, whereas PSNR and SSIM are measures of inverse distortion; this accounts for the negative slope in the first two figures, and the positive slope in the third.
-Figure 5 visually reinforces these findings by presenting selected recoveries ordered from low to high uncertainty.
-The results show a clear increase in perceptual quality with increasing uncertainty, accompanied by a visually subtle increase in distortion.
+Figure 6 displays similar trends for image inpainting, consistent with our super-resolution analysis and reinforcing the validity of our findings across diverse restoration tasks and data distributions.
+This is further visualized in Figure 2, which presents outputs from selected algorithms ordered by perceptual quality.
+The results clearly demonstrate an increase in hallucination (uncertainty) and distortion with increasing perceptual quality.
+Finally, Appendix H presents additional results obtained via direct estimation of statistics in high dimensions, further supporting our theoretical analysis.
+Figure 6: Evaluation of LDMs on image inpainting, highlighting the trade-off between uncertainty and perceptual quality (top) and the uncertainty-distortion relationship (bottom).
+No model achieves both low uncertainty and high perceptual quality, with higher uncertainty generally leading to increased distortion.
+Differing axis placements emphasize the distinct roles of uncertainty.
 6 Conclusion
-This study established the uncertainty-perception tradeoff in generative restoration, demonstrating that high perceptual quality leads to increased uncertainty, particularly in high dimensions.
+This study established the uncertainty-perception tradeoff in generative restoration, demonstrating that high perceptual quality leads to increased hallucination (uncertainty), particularly in high dimensions.
 We characterized this tradeoff and its fundamental relation to the inherent uncertainty of the problem, introducing the uncertainty-perception plane which may guide practitioners in understanding estimator performance.
@@
 7 Limitations
-While our theoretical analysis is applicable to any signal restoration task with non-invertible degradation, we focused our empirical validation on image super-resolution for consistency with prior art.
-Therefore, expanding the experiments to include additional image-to-image tasks and domains such as audio, video, and text could reveal broader implications and applications of this tradeoff.
 Our analysis is grounded in the theoretical framework of entropy as a measure of uncertainty.
-However, the practical estimation of high-dimensional entropy remains an active area of research.
-To address this, we utilize a tractable upper bound in Section 4.3.
-This alternative approach calls for further study, investigating its potential to effectively quantify uncertainty in analyzing the performance of existing algorithms.
-Furthermore, incorporating this upper bound into the design of new algorithms, whether as an objective function to minimize or a regularization term to constrain, could enable explicit control over the uncertainty-perception tradeoff, potentially leading to more reliable solutions.
+Information theory offers a powerful framework for quantifying uncertainty and dependencies in data, handling multivariate and heterogeneous data types, and capturing complex patterns.
+However, its wider adoption has been limited by the challenge of estimating information-theoretic measures in high dimensions.
+The curse of dimensionality makes accurate density estimation infeasible [12, 48], leading many to rely on simpler second-order statistics.
+The development of practical tools for estimating statistics in high-dimensional data remains an active area of research [76].
+While initial approaches assumed exponential family distributions (e.g., Gaussian) for tractable calculations [57], their performance degrades for long-tailed distributions.
+Non-parametric methods like binning strategies, including KDE and kNN estimators [61, 40, 29], offer more flexibility but are data-dependent and sensitive to parameter choices.
+Alternative approaches involve ensemble estimation [43] or von Mises Expansions [35], the distributional analog of the Taylor expansion.
+Rotation-Based Iterative Gaussianization [46] presents a promising direction by transforming data into a multivariate Gaussian domain, simplifying density estimation.
+However, its application to images has been limited to small patches due to the computational challenges of learning rotations based on principal or independent component analysis.
+A recent extension addresses this by utilizing convolutional rotations, enabling efficient processing of entire images [45].
+While accurately estimating high-dimensional entropy remains an active research area, Section 5 utilizes a tractable upper bound.
+This alternative calls for further investigation into its potential for quantifying uncertainty and analyzing algorithm performance.
+Moreover, incorporating this bound into the design of new algorithms could enable explicit control over the uncertainty-perception trade-off, potentially leading to more reliable solutions.
+Lastly, we focused our empirical validation on image super-resolution and inpainting, two benchmark problems in image restoration.
+Our analysis, however, applies to any restoration task with non-invertible degradation.
+Hence, expanding the experiments to additional image-to-image tasks and domains such as audio, video, and text may reveal broader implications and applications of our work.
 8 Broader Impact
 Our work revealing a fundamental tradeoff between uncertainty and perception in image restoration carries significant societal impact.
-This knowledge empowers users and developers to utilize and build generative models more responsibly, particularly in critical fields like healthcare, forensics, and autonomous systems.
-By understanding this inherent trade-off, practitioners can consider trading performance for better safety and resilience against potential misuse and misinterpretations.
-Appendix A Information-Theory Preliminaries
+Developers across various fields, including healthcare and autonomous systems, often integrate cutting-edge models into their applications, prioritizing state-of-the-art performance and perceptual quality.
+However, our work aims to highlight a crucial factor often overlooked: the inherent tradeoff between uncertainty and perception.
+By raising awareness of this tradeoff, we empower developers to make informed decisions that prioritize safety and reliability over purely perceptual enhancements.
+For instance, in healthcare, potential restoration algorithms can be evaluated by plotting them on the uncertainty-perception plane, facilitating the identification of methods that strike the optimal balance for specific clinical needs.
+Furthermore, by understanding this inherent trade-off, practitioners can consider trading performance for better safety and resilience against potential misuse and misinterpretations.
+While primarily theoretical, our analysis yields a practical measure of uncertainty (or entropy), used in our experiments to visually and quantitatively illustrate our findings.
+This tractable uncertainty measure, or any differentiable alternative, can be incorporated into a loss function during the training of generative models like GANs or as an optimization objective to guide the reverse process in diffusion models.
+This approach enables the development of algorithms that explicitly optimize for the tradeoff between uncertainty and perception.
+Appendix A Conditional Divergence and Human Perception
+In our context, perception is defined as the probability $p_{\text{success}}$ of a human observer successfully distinguishing between a pair of natural and degraded images, drawn from $p_{X,Y})$, and a pair of restored and degraded images drawn from $p_{\hat{X},Y})$.
+From a Bayesian perspective, the optimal decision rule maximizing $p_{\text{success}}$ yields ([56] Section 2):
+$p_{\text{success}}=\frac{1}{2}+\frac{1}{2}D_{\text{TV}}(p_{X,Y},p_{\hat{X},Y})$
+where $D_{\text{TV}}(p_{X,Y},p_{\hat{X},Y})$ is the total-variation (TV) distance.
+When $D(p_{X,Y},p_{\hat{X},Y})=0$, the two pairs are indistinguishable ($p_{\text{success}}=0.5$), implying perfect perception quality.
+We generalize this beyond the total-variation (TV) distance to any conditional divergence, recognizing that the divergence that best relates to human perception remains an open question.
+Appendix B Information-Theory Preliminaries
 To make the paper self-contained, we briefly overview the essential definitions and results in information-theory.
@@
 $h(X)+h(Y)=h(X_{g})+h(Y_{g})=h(X_{g}+Y_{g})\leq h(X+Y).$
-Appendix B Derivation of Example 1
+Appendix C Derivation of Example 1
 Since $\hat{X}=\mathbb{E}\left[X|Y\right]+Z$, then $\hat{X}|Y\sim{\mathcal{N}}(\mathbb{E}\left[X|Y\right],\sigma_{z}^{2})$.
@@
 Figure 7: The Uncertainty-Perception function for Example 1. As perception quality improves, the minimal achievable uncertainty increases, suggesting a tradeoff governed by the inherent uncertainty.
-Appendix C Proof of Theorem 1
+Appendix D Proof of Theorem 1
 First, the constraint ${\mathcal{C}}(P)\triangleq\{\hat{X}\,:\,D_{v}(X,\hat{X}\big{|}Y)\leq P$} defines a compact set which is continuous in $P$.
@@
 where the first inequality is due to Lemma 1. Finally, for any $P\geq 0$ it holds that $U(P)\leq U(0)$ which implies $U(0)\leq 2N(X_{G}|Y)$, completing the proof.
-Appendix D Proof of Theorem 2
+Appendix E Proof of Theorem 2
 Assuming $D_{v}(X,\hat{X}\big{|}Y)$ is convex in its second argument, the constraint represent a compact, convex set.
@@
 Therefore, we minimize a log-concave function over a convex domain and thus the global minimum is attained on the set boundary where $D_{v}(X,\hat{X}\big{|}Y)=P$.
-Appendix E Proof of Theorem 3
+Appendix F Proof of Theorem 3
 We begin with applying Lemma 1 and Lemma 2 to bound the objective function as follows
@@
 $U_{g}(P)=\eta(P)\cdot N(X|Y).$ (31)
-Appendix F Proof of Theorem 4
+Appendix G Proof of Theorem 4
 Define ${\mathcal{E}}\triangleq\hat{X}-X$. Then,
@@
 where (a) is by the law of total expectation, (b) is due to the inequality of arithmetic and geometric means, (c) follows Lemma 1, and (d) is according to Jensen’s inequality.
-Appendix G Results via Direct Estimation
+Appendix H Results via Direct Estimation
 Estimating high-dimensional statistics is prone to errors [46].
@@
 Results are consistent with the finding in Figure 5.
+NeurIPS Paper Checklist
+  1. 1.
+    Claims
+  2. Question: Do the main claims made in the abstract and introduction accurately reflect the paper’s contributions and scope?
+  3. Answer: [Yes]
+  4. Justification: We ensured the abstract and introduction the abstract and introduction describe our major contributions.
+  5. Guidelines:
+    • The answer NA means that the abstract and introduction do not include the claims made in the paper.
+    • The abstract and/or introduction should clearly state the claims made, including the contributions made in the paper and important assumptions and limitations.
+A No or NA answer to this question will not be perceived well by the reviewers.
+    • The claims made should match theoretical and experimental results, and reflect how much the results can be expected to generalize to other settings.
+    • It is fine to include aspirational goals as motivation as long as it is clear that these goals are not attained by the paper.
+  6. 2.
+    Limitations
+  7. Question: Does the paper discuss the limitations of the work performed by the authors?
+  8. Answer: [Yes]
+  9. Justification: Limitations of our work are discussed extensively in a dedicated section.
+  10. Guidelines:
+    • The answer NA means that the paper has no limitation while the answer No means that the paper has limitations, but those are not discussed in the paper.
+    • The authors are encouraged to create a separate "Limitations" section in their paper.
+    • The paper should point out any strong assumptions and how robust the results are to violations of these assumptions (e.g., independence assumptions, noiseless settings, model well-specification, asymptotic approximations only holding locally).
+The authors should reflect on how these assumptions might be violated in practice and what the implications would be.
+    • The authors should reflect on the scope of the claims made, e.g., if the approach was only tested on a few datasets or with a few runs.
+In general, empirical results often depend on implicit assumptions, which should be articulated.
+    • The authors should reflect on the factors that influence the performance of the approach.
+For example, a facial recognition algorithm may perform poorly when image resolution is low or images are taken in low lighting.
+Or a speech-to-text system might not be used reliably to provide closed captions for online lectures because it fails to handle technical jargon.
+    • The authors should discuss the computational efficiency of the proposed algorithms and how they scale with dataset size.
+    • If applicable, the authors should discuss possible limitations of their approach to address problems of privacy and fairness.
+    • While the authors might fear that complete honesty about limitations might be used by reviewers as grounds for rejection, a worse outcome might be that reviewers discover limitations that aren’t acknowledged in the paper.
+The authors should use their best judgment and recognize that individual actions in favor of transparency play an important role in developing norms that preserve the integrity of the community.
+Reviewers will be specifically instructed to not penalize honesty concerning limitations.
+  11. 3.
+    Theory Assumptions and Proofs
+  12. Question: For each theoretical result, does the paper provide the full set of assumptions and a complete (and correct) proof?
+  13. Answer: [Yes]
+  14. Justification: A detailed problem formulation, including all assumptions, is provided om a dedicated section. We have taken great care to ensure the clarity and correctness of our proofs.
+  15. Guidelines:
+    • The answer NA means that the paper does not include theoretical results.
+    • All the theorems, formulas, and proofs in the paper should be numbered and cross-referenced.
+    • All assumptions should be clearly stated or referenced in the statement of any theorems.
+    • The proofs can either appear in the main paper or the supplemental material, but if they appear in the supplemental material, the authors are encouraged to provide a short proof sketch to provide intuition.
+    • Inversely, any informal proof provided in the core of the paper should be complemented by formal proofs provided in appendix or supplemental material.
+    • Theorems and Lemmas that the proof relies upon should be properly referenced.
+  16. 4.
+    Experimental Result Reproducibility
+  17.
+Question: Does the paper fully disclose all the information needed to reproduce the main experimental results of the paper to the extent that it affects the main claims and/or conclusions of the paper (regardless of whether the code and data are provided or not)?
+  18. Answer: [Yes]
+  19.
+Justification: Although our main contribution is theoretical, we complement it with an empirical analysis of existing open models.
+This analysis is presented with complete details to ensure full reproducibility.
+  20. Guidelines:
+    • The answer NA means that the paper does not include experiments.
+    • If the paper includes experiments, a No answer to this question will not be perceived well by the reviewers: Making the paper reproducible is important, regardless of whether the code and data are provided or not.
+    • If the contribution is a dataset and/or model, the authors should describe the steps taken to make their results reproducible or verifiable.
+    • Depending on the contribution, reproducibility can be accomplished in various ways.
+For example, if the contribution is a novel architecture, describing the architecture fully might suffice, or if the contribution is a specific model and empirical evaluation, it may be necessary to either make it possible for others to replicate the model with the same dataset, or provide access to the model.
+In general.
+releasing code and data is often one good way to accomplish this, but reproducibility can also be provided via detailed instructions for how to replicate the results, access to a hosted model (e.g., in the case of a large language model), releasing of a model checkpoint, or other means that are appropriate to the research performed.
+    • While NeurIPS does not require releasing code, the conference does require all submissions to provide some reasonable avenue for reproducibility, which may depend on the nature of the contribution.
+For example
+      1. (a)
+        If the contribution is primarily a new algorithm, the paper should make it clear how to reproduce that algorithm.
+      2. (b)
+        If the contribution is primarily a new model architecture, the paper should describe the architecture clearly and fully.
+      3. (c)
+        If the contribution is a new model (e.g., a large language model), then there should either be a way to access this model for reproducing the results or a way to reproduce the model (e.g., with an open-source dataset or instructions for how to construct the dataset).
+      4. (d)
+        We recognize that reproducibility may be tricky in some cases, in which case authors are welcome to describe the particular way they provide for reproducibility.
+In the case of closed-source models, it may be that access to the model is limited in some way (e.g., to registered users), but it should be possible for other researchers to have some path to reproducing or verifying the results.
+  21. 5.
+    Open access to data and code
+  22. Question: Does the paper provide open access to the data and code, with sufficient instructions to faithfully reproduce the main experimental results, as described in supplemental material?
+  23. Answer: [No]
+  24.
+Justification: Our experimental analysis utilizes open-source models and datasets.
+While our current submission does not include the code, we provide a comprehensive description of our experimental setup to facilitate reproduction of the results.
+  25. Guidelines:
+    • The answer NA means that paper does not include experiments requiring code.
+    • Please see the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
+    • While we encourage the release of code and data, we understand that this might not be possible, so “No” is an acceptable answer.
+Papers cannot be rejected simply for not including code, unless this is central to the contribution (e.g., for a new open-source benchmark).
+    • The instructions should contain the exact command and environment needed to run to reproduce the results.
+See the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
+    • The authors should provide instructions on data access and preparation, including how to access the raw data, preprocessed data, intermediate data, and generated data, etc.
+    • The authors should provide scripts to reproduce all experimental results for the new proposed method and baselines.
+If only a subset of experiments are reproducible, they should state which ones are omitted from the script and why.
+    • At submission time, to preserve anonymity, the authors should release anonymized versions (if applicable).
+    • Providing as much information as possible in supplemental material (appended to the paper) is recommended, but including URLs to data and code is permitted.
+  26. 6.
+    Experimental Setting/Details
+  27. Question: Does the paper specify all the training and test details (e.g., data splits, hyperparameters, how they were chosen, type of optimizer, etc.) necessary to understand the results?
+  28. Answer: [Yes]
+  29. Justification: Our analysis centers on pre-trained models applied to open datasets. Section 5 and Appendix H provide the technical details necessary for reproducing our results.
+  30. Guidelines:
+    • The answer NA means that the paper does not include experiments.
+    • The experimental setting should be presented in the core of the paper to a level of detail that is necessary to appreciate the results and make sense of them.
+    • The full details can be provided either with the code, in appendix, or as supplemental material.
+  31. 7.
+    Experiment Statistical Significance
+  32. Question: Does the paper report error bars suitably and correctly defined or other appropriate information about the statistical significance of the experiments?
+  33. Answer: [No]
+  34.
+Justification: Our experimental analysis involves applying pre-trained models to a fixed set of open datasets, resulting in deterministic outputs.
+As there is no inherent randomness or variation in the experimental process, traditional statistical significance measures like error bars or confidence intervals are not applicable.
+  35. Guidelines:
+    • The answer NA means that the paper does not include experiments.
+    • The authors should answer "Yes" if the results are accompanied by error bars, confidence intervals, or statistical significance tests, at least for the experiments that support the main claims of the paper.
+    • The factors of variability that the error bars are capturing should be clearly stated (for example, train/test split, initialization, random drawing of some parameter, or overall run with given experimental conditions).
+    • The method for calculating the error bars should be explained (closed form formula, call to a library function, bootstrap, etc.)
+    • The assumptions made should be given (e.g., Normally distributed errors).
+    • It should be clear whether the error bar is the standard deviation or the standard error of the mean.
+    • It is OK to report 1-sigma error bars, but one should state it.
+The authors should preferably report a 2-sigma error bar than state that they have a 96% CI, if the hypothesis of Normality of errors is not verified.
+    • For asymmetric distributions, the authors should be careful not to show in tables or figures symmetric error bars that would yield results that are out of range (e.g. negative error rates).
+    • If error bars are reported in tables or plots, The authors should explain in the text how they were calculated and reference the corresponding figures or tables in the text.
+  36. 8.
+    Experiments Compute Resources
+  37. Question: For each experiment, does the paper provide sufficient information on the computer resources (type of compute workers, memory, time of execution) needed to reproduce the experiments?
+  38. Answer: [No]
+  39.
+Justification: Our primary contribution is theoretical, and the accompanying experimental analysis is computationally lightweight, requiring only basic processing on a standard CPU given the ground-truth, distorted, and recovered images.
+Therefore, detailed compute resource specifications are not essential for reproducing the results.
+  40. Guidelines:
+    • The answer NA means that the paper does not include experiments.
+    • The paper should indicate the type of compute workers CPU or GPU, internal cluster, or cloud provider, including relevant memory and storage.
+    • The paper should provide the amount of compute required for each of the individual experimental runs as well as estimate the total compute.
+    • The paper should disclose whether the full research project required more compute than the experiments reported in the paper (e.g., preliminary or failed experiments that didn’t make it into the paper).
+  41. 9.
+    Code Of Ethics
+  42. Question: Does the research conducted in the paper conform, in every respect, with the NeurIPS Code of Ethics https://neurips.cc/public/EthicsGuidelines?
+  43. Answer: [Yes]
+  44. Justification: We confirm that our study aligns with the NeurIPS Code of Ethics.
+  45. Guidelines:
+    • The answer NA means that the authors have not reviewed the NeurIPS Code of Ethics.
+    • If the authors answer No, they should explain the special circumstances that require a deviation from the Code of Ethics.
+    • The authors should make sure to preserve anonymity (e.g., if there is a special consideration due to laws or regulations in their jurisdiction).
+  46. 10.
+    Broader Impacts
+  47. Question: Does the paper discuss both potential positive societal impacts and negative societal impacts of the work performed?
+  48. Answer: [Yes]
+  49. Justification: Broader Impacts of our work are discussed in their own dedicated section.
+  50. Guidelines:
+    • The answer NA means that there is no societal impact of the work performed.
+    • If the authors answer NA or No, they should explain why their work has no societal impact or why the paper does not address societal impact.
+    • Examples of negative societal impacts include potential malicious or unintended uses (e.g., disinformation, generating fake profiles, surveillance), fairness considerations (e.g., deployment of technologies that could make decisions that unfairly impact specific groups), privacy considerations, and security considerations.
+    • The conference expects that many papers will be foundational research and not tied to particular applications, let alone deployments.
+However, if there is a direct path to any negative applications, the authors should point it out.
+For example, it is legitimate to point out that an improvement in the quality of generative models could be used to generate deepfakes for disinformation.
+On the other hand, it is not needed to point out that a generic algorithm for optimizing neural networks could enable people to train models that generate Deepfakes faster.
+    • The authors should consider possible harms that could arise when the technology is being used as intended and functioning correctly, harms that could arise when the technology is being used as intended but gives incorrect results, and harms following from (intentional or unintentional) misuse of the technology.
+    • If there are negative societal impacts, the authors could also discuss possible mitigation strategies (e.g., gated release of models, providing defenses in addition to attacks, mechanisms for monitoring misuse, mechanisms to monitor how a system learns from feedback over time, improving the efficiency and accessibility of ML).
+  51. 11.
+    Safeguards
+  52.
+Question: Does the paper describe safeguards that have been put in place for responsible release of data or models that have a high risk for misuse (e.g., pretrained language models, image generators, or scraped datasets)?
+  53. Answer: [N/A]
+  54. Justification: This paper focuses on analyzing existing open-source models and datasets, and therefore does not introduce new models or datasets that require specific safeguards.
+  55. Guidelines:
+    • The answer NA means that the paper poses no such risks.
+    • Released models that have a high risk for misuse or dual-use should be released with necessary safeguards to allow for controlled use of the model, for example by requiring that users adhere to usage guidelines or restrictions to access the model or implementing safety filters.
+    • Datasets that have been scraped from the Internet could pose safety risks. The authors should describe how they avoided releasing unsafe images.
+    • We recognize that providing effective safeguards is challenging, and many papers do not require this, but we encourage authors to take this into account and make a best faith effort.
+  56. 12.
+    Licenses for existing assets
+  57.
+Question: Are the creators or original owners of assets (e.g., code, data, models), used in the paper, properly credited and are the license and terms of use explicitly mentioned and properly respected?
+  58. Answer: [Yes]
+  59. Justification: All relevant publicly-available models and datasets utilized in our work are properly cited and acknowledged in the paper.
+  60. Guidelines:
+    • The answer NA means that the paper does not use existing assets.
+    • The authors should cite the original paper that produced the code package or dataset.
+    • The authors should state which version of the asset is used and, if possible, include a URL.
+    • The name of the license (e.g., CC-BY 4.0) should be included for each asset.
+    • For scraped data from a particular source (e.g., website), the copyright and terms of service of that source should be provided.
+    • If assets are released, the license, copyright information, and terms of use in the package should be provided.
+For popular datasets, paperswithcode.com/datasets has curated licenses for some datasets.
+Their licensing guide can help determine the license of a dataset.
+    • For existing datasets that are re-packaged, both the original license and the license of the derived asset (if it has changed) should be provided.
+    • If this information is not available online, the authors are encouraged to reach out to the asset’s creators.
+  61. 13.
+    New Assets
+  62. Question: Are new assets introduced in the paper well documented and is the documentation provided alongside the assets?
+  63. Answer: [N/A]
+  64. Justification: We do not release new assets.
+  65. Guidelines:
+    • The answer NA means that the paper does not release new assets.
+    • Researchers should communicate the details of the dataset/code/model as part of their submissions via structured templates. This includes details about training, license, limitations, etc.
+    • The paper should discuss whether and how consent was obtained from people whose asset is used.
+    • At submission time, remember to anonymize your assets (if applicable). You can either create an anonymized URL or include an anonymized zip file.
+  66. 14.
+    Crowdsourcing and Research with Human Subjects
+  67.
+Question: For crowdsourcing experiments and research with human subjects, does the paper include the full text of instructions given to participants and screenshots, if applicable, as well as details about compensation (if any)?
+  68. Answer: [N/A]
+  69. Justification: The paper does not involve crowdsourcing nor research with human subjects.
+  70. Guidelines:
+    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
+    • Including this information in the supplemental material is fine, but if the main contribution of the paper involves human subjects, then as much detail as possible should be included in the main paper.
+    • According to the NeurIPS Code of Ethics, workers involved in data collection, curation, or other labor should be paid at least the minimum wage in the country of the data collector.
+  71. 15.
+    Institutional Review Board (IRB) Approvals or Equivalent for Research with Human Subjects
+  72.
+Question: Does the paper describe potential risks incurred by study participants, whether such risks were disclosed to the subjects, and whether Institutional Review Board (IRB) approvals (or an equivalent approval/review based on the requirements of your country or institution) were obtained?
+  73. Answer: [N/A]
+  74. Justification: The paper does not involve crowdsourcing nor research with human subjects.
+  75. Guidelines:
+    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
+    • Depending on the country in which research is conducted, IRB approval (or equivalent) may be required for any human subjects research.
+If you obtained IRB approval, you should clearly state this in the paper.
+    • We recognize that the procedures for this may vary significantly between institutions and locations, and we expect authors to adhere to the NeurIPS Code of Ethics and the guidelines for their institution.
+    • For initial submissions, do not include any information that would break anonymity (if applicable), such as the institution conducting the review.